OpenStudy (anonymous):

What is the sum of the geometric sequence 8, -16, 32 . if there are 15 terms?

OpenStudy (anonymous):

a=8 r=-16/8=-2 n=15 \[S _{n}=\frac{ a \left( r^n-1 \right) }{ r-1 }\]

OpenStudy (anonymous):

Thanks so much~

OpenStudy (anonymous):

yw